Anggy:

A triangle has side lengths of 8cm 15cm and 26cm classify it as acute obtuse or right

MrMudd183:

What do you think the first thing to do is?

luvvdani1516:

An easy way to figure this out is using the Pythagorean Theorem. Or \[a ^{2}+b ^{2}=c ^{2}\]

mhanifa:

Use Pythagorean and find if c^2 = < or > compared to a^2 + b^2

luvvdani1516:

@luvvdani1516 wrote:
An easy way to figure this out is using the Pythagorean Theorem. Or \[a ^{2}+b ^{2}=c ^{2}\]
C is always the longest side of the triangle. Start with that, if you need more explaining, I'm here to help.

Extrinix:

Right: \(c^2=a^2+b^2\) Obtuse: \(c^2 \gt a^2+b^2\) Acute: \(c^2 \lt a^2+b^2\) Here, this should help.

mhanifa:

When you get = you know the triangle is right

mhanifa:

c^2 > a^2 + b^2 means the angle opposite to c is greater than 90 deg, hence the triangle is obtuse